# Cosmology with non-conformal holographic matter

### v1

@@ 1 Introduction
There have been many other works in the past investigating the thermodynamic behavior of holographic field theories or aspects of their cosmological physics. Papers that specifically consider aspects of equilibrium and non-equilibrium thermodynamics for non-conformal holographic theories include Attems_2016 ; G_rsoy_2018 ; Kleinertt_2016 ; Attems_2017 ; Ballon_QCD_2021 ; Buchel_QCD_2016 ; G_rsoy_QCD_2016 . Various papers have studied the behavior of holographic field theories on fixed cosmological backgrounds, including Ghosh_2018 ; Casalderrey_Solana_2021 ; Marolf_2011 . Other papers that have studied cosmological evolution of holographic field theories dynamically coupled to gravity include 2022paper ; dynamicalinflaton ; Koyama2001 where the authors go beyond the quasi-static approximation to study situations where the quantum field theory may be out of equilibrium. In reheating ; reheating2 ; reheating3 ; warminflation the authors consider holographic field theories as part of the physics of inflation; an important aspect of strongly coupled QFTs is that they quickly thermalize and enter the hydyodynamic regime making them potentially useful for facilitating the exit from inflation to the hot Big Bang. Novel aspects of our paper compared to previous works include i) a detailed investigation of how the scale dependent equation of state parameter depends on parameters (dimension of perturbing operator dimension and length of RG flow) in perhaps the simplest possible example of a holographic field theory with an RG flow between UV and IR CFT ii) a different numerical method that does not make use of the counterterms in the gravitational action and iii) a numerical study of the dynamical evolution of the scale factor for flat cosmologies where the only matter is the non-conformal holographic theory.

### v2

@@ 1 Introduction
There have been many other works in the past investigating the thermodynamic behavior of holographic field theories or aspects of their cosmological physics. Papers that specifically consider aspects of equilibrium and non-equilibrium thermodynamics for non-conformal holographic theories include Attems_2016 ; G_rsoy_2018 ; Kleinertt_2016 ; Attems_2017 ; Ballon_QCD_2021 ; Buchel_QCD_2016 ; G_rsoy_QCD_2016 . Various papers have studied the behavior of holographic field theories on fixed cosmological backgrounds, including Ghosh_2018 ; Casalderrey_Solana_2021 ; Marolf_2011 . Other papers that have studied cosmological evolution of holographic field theories dynamically coupled to gravity include 2022paper ; dynamicalinflaton ; Koyama2001 where the authors go beyond the quasi-static approximation to study situations where the quantum field theory may be out of equilibrium. Holographic non-conformal matter has also been studied previously in the context of braneworld cosmology, with equation-of-state parameters arising from scalar potentials different from those considered in this paper (for related work, see Fichet:2022ixi ; Fichet:2022xol ; Fichet:2023dju ). In reheating ; reheating2 ; reheating3 ; warminflation the authors consider holographic field theories as part of the physics of inflation; an important aspect of strongly coupled QFTs is that they quickly thermalize and enter the hydyodynamic regime making them potentially useful for facilitating the exit from inflation to the hot Big Bang. Novel aspects of our paper compared to previous works include i) a detailed investigation of how the scale dependent equation of state parameter depends on parameters (dimension of perturbing operator and length of RG flow) in perhaps the simplest possible example of a holographic field theory with an RG flow between UV and IR CFT ii) a different numerical method that does not make use of the counterterms in the gravitational action and iii) a numerical study of the dynamical evolution of the scale factor for flat cosmologies where the only matter is the non-conformal holographic theory.